Post-laboratory Questions: Week 8, Determination of Ksp 1. Using your calculated value of Kap, determine the molar solubility of Agzcro 2. You determined the concentration of the Cros2-spectrophotometrically, and used that concentration to help determine Ksp. What would happen to the value of Kp (higher, lower, or no change) ere was an error in your Spec20 and all of your measured values for cro42 wereo.12 AU too large? Why?
media%2F77f%2F77f0e0d8-10ef-4862-9c63-36
media%2F252%2F252a3130-f7fa-4d7f-97ca-eb
media%2Fc6f%2Fc6fbfd8a-1369-477d-87db-3d
Answer #1

Q1.

Ksp = 9*10^-12 approx

so

Ag2CrO4 <-> 2Ag+ + CrO4-2

Ksp = [Ag+][CrO4-2]

let 1 mol of salt be "S"

so

[Ag+] = 2S, CrO4-2 = S

Ksp = (2S)^2(S)

9*10^-12 = 4*S^3

S = ((9*10^-12)/(4)) ^(1/3)

S = 1.3103*10^-4